OpenStudy (anonymous):

The sum of two numbers is 60. The greater number is 6 more than the smaller number. Which equation can be used to solve for the smaller number? x + (x + 6) = 60 x − (x + 6) = 60 x(x + 6) = 60 x(x− 6) = 60

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

small number = x big number = (x + 6) small number + big number = 60

OpenStudy (anonymous):

X + ( x+6) @Vocaloid

Vocaloid (vocaloid):

yes

OpenStudy (anonymous):

A

Vocaloid (vocaloid):

great

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

Brenda needs $40 to buy a large box of chocolate. She has saved $12 and plans to work as a babysitter to earn $7 per hour. Which inequality shows the minimum number of hours, n, that Brenda should work as a babysitter to earn enough to buy the large box of chocolate? 12 + 7n ≤ 40, so n ≤ 4 7n ≤ 40 + 12, so n ≤ 7.4 12 + 7n ≥ 40, so n ≥ 4 7n ≥ 40 + 12, so n ≥ 7.4

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

amount saved + (hourly wage)(number of hours) >= price of chocolate

OpenStudy (anonymous):

c?

Vocaloid (vocaloid):

right

OpenStudy (anonymous):

Martha works in a toy shop and earns $44 per day. She earns an extra $2 for each toy she sells. If Martha wants to earn at least $80 per day, which inequality shows the minimum number of toys, n, that she should sell? 44 + 2n ≥ 80, so n ≥ 18 44 + 2n ≤ 80, so n ≤ 18 44 + 2n ≥ 80, so n ≥ 34 44 + 2n ≤ 80, so n ≤ 34

Vocaloid (vocaloid):

same rule as the last one

OpenStudy (anonymous):

A?

Vocaloid (vocaloid):

yes

OpenStudy (anonymous):

Neil wants to buy some video games over the Internet. Each game costs $25.01 and has a shipping cost of $9.98 per order. If Neil wants to spend no more than $60 for his video games, which inequality shows the maximum number of video games, p, that he can buy? 9.98 + 25.01p ≤ 60, so p ≤ 2 9.98 − 25.01p ≤ 60, so p ≤ 4 9.98p + 25.01p ≤ 60, so p ≤ 5 9.98p − 25.01p ≤ 60, so p ≤ 6

Vocaloid (vocaloid):

(shipping price)(number of games) + (game cost)(number of games) <= 60

OpenStudy (anonymous):

c?

Vocaloid (vocaloid):

nope solve for p 9.98 + 25.01p ≤ 60

OpenStudy (anonymous):

A?

Vocaloid (vocaloid):

right

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

perimeter means add the sides togther

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

yes